OpenStudy (anonymous):

Find x and v in terms of c and b: (x/c)+(v/b)=1 (x/b)+(v/c)=1 x=? v=?

OpenStudy (anonymous):

So what exactly is this answer for this? I'm still confused with this.

OpenStudy (anonymous):

just with pencil and paper i get \[x=\frac{bc}{b+c}\]

OpenStudy (anonymous):

i will write the steps but let me check it first

OpenStudy (anonymous):

Isolate v first then plug it into the second equation

OpenStudy (anonymous):

yeah looks good to me

OpenStudy (anonymous):

\[v = b(1-\frac{x}{c})\]

OpenStudy (anonymous):

\[\frac{x}{b}+\frac{b(1-\frac{x}{c})}{c} = 1\]

OpenStudy (anonymous):

\[xc +b^{2}(1-\frac{x}{c})=bc\]

OpenStudy (anonymous):

i did this \[bx+cv=bc\] \[cx+bv=bc\] \[b^2x+bcv=b^2c\] \[c^2x+bc^2=bc^2\] subtract to get \[(b^2-c^2)x=b^2c-bc^2=bc(b-c)\] \[x=\frac{bc(b-c)}{b^2-c^2}=\frac{bc(b-c)}{(b-c)(b+c)}=\frac{bc}{b-c}\]

OpenStudy (anonymous):

you can also use cramer's rule i guess. oh and since this is symmetric in x and v, you get the same thing for v

OpenStudy (anonymous):

Shouldn't it be b+c ?